# UPGS: Unified Pose‑aware Gaussian Splatting for Dynamic Scene Deblurring

### v2

@@ 4.1 Implementation details
Dataset. We evaluate our method on the Stereo Blur dataset [34], which comprises six scenes exhibiting significant motion blur from both camera and object motion. Each sequence was captured with a ZED stereo camera: the left view provides the blurry input, while the right view serves as the sharp ground truth. Details of the blur synthesis process are described in [49]. For each scene, we extract 24 frames, and obtain camera extrinsics using COLMAP [32].
@@
Metrics. We perform thorough quantitative comparisons by measuring the synthesized novel views with established metrics: Peak Signal-to-Noise Ratio (PSNR) [5], Structural Similarity Index Measure (SSIM) [40], and Learned Perceptual Image Patch Similarity (LPIPS) [46], which together capture both reconstruction fidelity and perceptual quality.
Deblur4DGS Ours
PSNR / SSIM / LPIPS PSNR / SSIM / LPIPS
Street 29.508 / 0.905 / 0.137 30.383 / 0.920 / 0.096
Skating 30.893 / 0.900 / 0.158 31.877 / 0.915 / 0.122
Seesaw 29.777 / 0.092 / 0.101 29.786 / 0.929 / 0.096
Man 25.602 / 0.825 / 0.225 28.766 / 0.906 / 0.092
Women 25.789 / 0.806 / 0.242 27.693 / 0.869 / 0.137
Third 31.051 / 0.915 / 0.155 32.352 / 0.932 / 0.101
Table 2: Per-scene results on the Stereo Blur dataset compared to Deblur4DGS [42].
@@ 4.2 Comparison with State-of-the-Art Methods
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Affine 29.306 0.893 0.123
w/o Stagewise 28.778 0.879 0.168
Ours 30.143 0.911 0.107
@@
Quantitative Results. The quantitative results on the Stereo Blur dataset [34] are summarized in Table 1. We compare our pipeline among DyBluRF [34], Shape-of-Motion [37], 4DGS [41], and Deblur4DGS [42]. Our method attains the best PSNR and lowest LPIPS among all baselines, and a high SSIM that is close to the top score. Among the most comparable baselines, Deblur4DGS [42] shares a Gaussian-splatting backbone and is designed specifically for motion-blurred novel-view synthesis. Against it, we improve PSNR by 1.12 dB, SSIM by 0.166, and reduce LPIPS by 0.058 ($\approx$35% relative). These gains indicate lower reconstruction error, higher structural fidelity, and substantially better perceptual quality, corresponding to sharper textures, cleaner boundaries, and fewer blur artifacts in the renders. In addition, as illustrated in Table 2, our method achieves more significant improvements on scenes with challenging blurs, such as street, skating, and women. This further validates the superiority of our method.
@@ 4.3.1 Affine Transformations
Table 4: Effect of the number of frames on reconstruction quality and training time. Red and yellow denote the best and second best performances respectively.
@@ 4.3.3 Latent Sharp Frame NN
We also conduct experiments on $N$ to investigate the number of latent sharp frames within an exposure, aiming to balance reconstruction quality and training time. As stated in Sec 3.1, $N$ is the count of latent sharp frames whose temporal average reproduces the observed blurry image. We sweep $N=7$ to $13$. Table 4 shows clear gains up to $N=7$ and only marginal improvement beyond, while computational cost grows roughly linearly. We therefore set $N=7$ for our final configuration.
@@ 5 Conclusion
Supplementary Material for
UPGS: Unified Pose-aware Gaussian Splatting for Dynamic Scene Deblurring
6 Unified Pose-aware Formulation
@@
As camera poses are not differentiable in the 3DGS framework, previous methods commonly calculate camera poses before optimizing Gaussians without fully considering their coupled relationship. As a result, these methods suffer inferior performance when blurry artifacts undermine the pose estimation accuracy. With our pose-aware formulation, we can introduce camera pose as learnable affine transformations that are compatible with 3DGS framework for end-to-end optimization, thereby producing superior performance.
7 Additional Ablation Results
The ablation results on affine transformations and our training strategy across all six scenes are illustrated in Table 5, 6, 7, which further demonstrate the effectiveness of our method designs.
8 Demo videos
We also provide demo videos for qualitative evaluation under real-world conditions (please see the “demo” folder in the supplementary material). Compared with the baseline, our renderings demonstrate noticeably improved temporal consistency: object contours remain stable without jitter, textures are coherent over time with minimal flicker or shimmer, and motion-blur streaks appear smooth and continuous. These improvements are particularly evident in challenging regions where fast object motion overlaps with strong camera shake. The supplementary demo footage provides visual evidence of these effects, showcasing sharper, more stable novel views that complement the quantitative results presented in the paper.
Figure 6:
@@ 8 Demo videos
Table 7: LPIPS of the ablation studies on all scene

### v3

@@ 4.1 Implementation details
Dataset. We evaluate our method on the Stereo Blur dataset [sun2024dyblurf] and BARD-GS dataset [lu2025bard]. Stereo Blur dataset comprises six scenes exhibiting significant motion blur from both camera and object motion. Each sequence was captured with a ZED stereo camera: the left view provides the blurry input, while the right view serves as the sharp ground truth. Details of the blur synthesis process are described in [zhou2019davanet]. For each scene, we extract 24 frames, and obtain camera extrinsics using COLMAP [schonberger2016structure]. BARD-GS dataset are specifically collected for dynamic deblurring task. We further subselect frames that exhibit the most representative motions within each scene.
@@
Metrics. We perform thorough quantitative comparisons by measuring the synthesized novel views with established metrics: Peak Signal-to-Noise Ratio (PSNR) [huynh2008scope], Structural Similarity Index Measure (SSIM) [wang2004image], and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], which together capture both reconstruction fidelity and perceptual quality.
@@ 4.2 Comparison with State-of-the-Art Methods
Table 2: Deblurring results across three regions on the Stereo Blur dataset
Methods Dynamic region Static region Edge region
[…]
Deblur4DGS 28.283 / 0.817 / 0.125 29.672 / 0.716 / 0.182 24.844 / 0.869 / 0.135
Ours 29.315 / 0.955 / 0.056 30.531 / 0.878 / 0.158 26.511 / 0.988 / 0.021
@@
Quantitative Results. The quantitative results on the Stereo Blur [sun2024dyblurf] and BARD-GS dataset [lu2025bard] are summarized in Table 1. We compare our pipeline among DyBluRF [sun2024dyblurf], Shape-of-Motion [wang2024shape], 4DGS [wu20244d], and Deblur4DGS [wu2024deblur4dgs]. Our method attains the best PSNR and lowest LPIPS among all baselines, and a high SSIM that is close to the top score. Among the most comparable baselines, Deblur4DGS [wu2024deblur4dgs] shares a Gaussian-splatting backbone and is designed specifically for motion-blurred novel-view synthesis. Against it, we improve PSNR by 1.12 dB, SSIM by 0.166, and reduce LPIPS by 0.058 ($\approx$35% relative). These gains indicate lower reconstruction error, higher structural fidelity, and substantially better perceptual quality, corresponding to sharper textures, cleaner boundaries, and fewer blur artifacts in the renders. In addition, Table 2 reports the metrics on three regions—dynamic (FG), static (BG), and edges—to pinpoint where deblurring quality changes. The edge region is a boundary band around moving objects, where pose and motion errors couple most and artifacts tend to concentrate. Compared with Deblur4DGS (the closest 4DGS deblurring baseline), our method improves all three regions, with the largest gains on edges. As illustrated in the supplementary, our method achieves more significant improvements on scenes with challenging blurs, such as street, skating, and women. This further validates the superiority of our method.
@@ 4.3.1 Affine Transformations
Stereo Dataset BARD-GS Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Affine 29.306 0.893 0.123 18.772 0.734 0.528
w/o Stagewise 28.778 0.879 0.168 18.421 0.677 0.531
Ours 30.143 0.911 0.107 19.502 0.731 0.550
Table 4: Ablations on number of latent frame and reference-frame evaluation. Red and Yellow indicate best and second best.
@@
(a) Effect of frame count.
Reference Frame Stereo Blur BARD
[…]
Middle 29.17 0.867 0.134 18.33 0.680 0.624
Sharpest (FG) 29.49 0.874 0.114 19.19 0.725 0.596
Sharpest 29.72 0.898 0.111 19.08 0.722 0.601
Ours 30.14 0.911 0.107 19.50 0.731 0.550
(b) Reference pose comparison.
@@ 4.3.3 Latent Sharp Frame NN
We also conduct experiments on $N$ in Stereo Blur dataset, to investigate the number of latent sharp frames within an exposure, aiming to balance reconstruction quality and training time. As stated in Sec 3.1, $N$ is the count of latent sharp frames whose temporal average reproduces the observed blurry image. We sweep $N=7$ to $13$. Table 4(a) shows clear gains up to $N=7$ and only marginal improvement beyond, while computational cost grows roughly linearly. We therefore set $N=7$ for our final configuration.
4.3.4 Reference pose sensitivity
We compare different reference pose selections in Table 4(b), including the middle subframe, the sharpest subframe in the foreground region, and the sharpest subframe over the entire image. In our method, we select as reference the subframe that contains the largest number of visible foreground Gaussians. Empirically, this choice yields consistently strong performance across datasets. While selecting the sharpest foreground subframe produces comparable results, the performance differences are generally small. In our experiments, the sharpest subframe often lies close to the subframe with maximal foreground visibility, suggesting that these choices lead to similar optimization behavior in this setting.
@@ 5 Conclusion
Supplementary Material
Appendix 0.A Unified Pose-aware Formulation
@@
In the presence of motion blur, pose inaccuracies are prone to be absorbed into the scene geometry. Previous methods commonly use image-loss to implicitly drive camera pose optimization. As a result, these methods suffer inferior performance when blurry artifacts undermine the pose estimation accuracy. With our pose-aware formulation, we can introduce camera pose as learnable affine transformations that are compatible with 3DGS framework for end-to-end optimization, thereby producing superior performance.
Appendix 0.B Gating Hyperparameter Specification
To ensure reproducible gating behaviour across all experiments, we report the exact hyperparameter values used for the two adaptive gates as shown in Table 5.
Stereo Blur BARD-GS
$\epsilon_{1}$ $2.4\times 10^{-2}$ $4.0\times 10^{-2}$
$\tau_{a}$ $3.92\times 10^{-5}$ $1.92\times 10^{-5}$
$\epsilon_{2}$ $7.0\times 10^{-3}$ $4.46\times 10^{-5}$
$\tau_{\mathrm{gain}}$ $7.3\times 10^{-4}$ $3.0\times 10^{-3}$
Table 5: Gating hyperparameters for the two adaptive gating criteria
Appendix 0.C Additional Results and Ablations
We record the full quantitative results for all scenes across both datasets, comparing our method against existing baselines. The metrics are summarized in Table 6. The ablation results on affine transformations and our training strategy across all six scenes are illustrated in Table 7, which further demonstrate the effectiveness of our method designs.
Card Toycar Poster
Method PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
4DGS 14.04 0.54 0.99 14.98 0.56 0.91 20.10 0.70 0.80
Shape-of-motion 15.41 0.74 0.50 16.25 0.75 0.43 20.86 0.75 0.40
Deblur4DGS 18.15 0.75 0.32 16.63 0.77 0.38 21.72 0.74 0.33
Ours 18.05 0.77 0.53 18.72 0.80 0.55 22.80 0.76 0.52
Windmill Kitchen Shark-spin
Method PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
4DGS 18.21 0.69 0.81 17.70 0.54 0.91 18.66 0.63 0.83
Shape-of-motion 17.75 0.76 0.41 18.59 0.62 0.45 18.65 0.68 0.42
Deblur4DGS 17.99 0.74 0.39 18.61 0.64 0.42 19.71 0.70 0.39
Ours 18.08 0.78 0.57 19.25 0.61 0.58 20.12 0.68 0.55
Table 6: Quantitative comparison across all six scenes (Card, Toycar, Poster, Windmill, Kitchen, Shark-spin) on the BARD-GS dataset. Metrics include PSNR, SSIM, and LPIPS.
Appendix 0.D Demo videos
We also provide demo.mov for qualitative comparisons under real-world conditions. Compared with the baseline, our renderings demonstrate noticeably improved temporal consistency: object contours remain stable without jitter, textures are coherent over time with minimal flicker or shimmer, and motion-blur streaks appear smooth and continuous. These improvements are particularly evident in challenging regions where fast object motion overlaps with strong camera shake. The supplementary demo footage provides visual evidence of these effects, showcasing sharper, more stable novel views that complement the quantitative results presented in the paper.
Figure 6: Validation of the unified pose-aware formulation. Rendering the transformed Gaussians from a fixed reference pose produces images that are nearly identical to those rendered directly from the source poses, and the error maps confirm only negligible differences.
@@ Appendix 0.D Demo videos
Table 7: Quantitative comparison across all six scenes (Street, Skating, Seesaw, Man, Women, Third) on the Stereo Blur dataset. Metrics include PSNR, SSIM, and LPIPS.